OpenStudy (anonymous):

Let f(x)= x-3 and g(x)= 2x+5. What is h(x) if f(h(x))= √x?

OpenStudy (anonymous):

Alright~ f(x)=x-3 therefore, f(h(x)) can also be expressed as (h(x))-3 = \[\sqrt{x}\] Then solve h(x)= √x +3

OpenStudy (anonymous):

\[f(h(x)) = h(x) -3 = \sqrt{x} \Rightarrow h(x) = \sqrt{x} +3\]